OpenStudy (anonymous):

What is the solution to the equation?-7(x-4)+8x=12+14

OpenStudy (anonymous):

@cwrw238

OpenStudy (igreen):

\(-7(x-4)+8x=12+14\) Distribute the -7 into the parenthesis: \(-7x+28+8x=12+14\) Add: \(-7x+28+8x=26\) Subtract 28 to both sides: \(-7x+8x=-2\) Combine like terms: \(x = -2\)

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @offsettakkoff

OpenStudy (igreen):

\(\bbox[5pt, #000000 ,border:5px solid #cbc14f]{\bf\huge \color{#cbc14f}\sim \! \! \! \! \color{#cbc14f}i\it\Huge\color{#cbc14f}G\color{#cbc14f}r\color{#cbc14f}e\color{#cbc14f}e\color{#cbc14f}n\bf\color{#cbc14f}ツ}\)